1) Find the maximum revenue if the price of selling x shoes is p(x) = 2 - 3x
Answer #1

solution: , given P(x) = 2-3x then R(x) P(x) .x ki = (2-3xx 2x + 3x² for maximum Revenue d R(2) =0 dx d {2x-3x²} - 0 da 2-61-